OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 5 - 15 + 45 - 135 + ... HELPPPP

OpenStudy (tkhunny):

Do those terms look like an arithmetic sequence? Geometric?

OpenStudy (anonymous):

geometric

OpenStudy (anonymous):

\[\sum_{n=0}^{infinity}5(-3)^n\] is this the answer?

OpenStudy (tkhunny):

That was quite a leap. Where did you get that?

OpenStudy (anonymous):

Well yes that would be :) I can just edit this a little for you :) \[\sum^\infty_{n=0}{5(-3)^n}\]

OpenStudy (anonymous):

I got that because i know that the lower limit is 0 but im not sure of the index...I do know that the ratio is -3 and the first term is 5

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

@KeithAfasCalcLover can you help really fast with another?

OpenStudy (anonymous):

Yeah sure!

OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 64 + 81 + 100 + 121 + ... + n2 + ... would it be \[\sum_{n=8}^{\infty}n^2\]

OpenStudy (tkhunny):

Well, good work, then. You seem to have reasoned it out ofter realizing that it was geometric.

OpenStudy (anonymous):

thanks :)

OpenStudy (tkhunny):

The last one looks fine. Neither geometric nor arithmetic.

OpenStudy (anonymous):

thank you!

OpenStudy (tkhunny):

...and you found the \(\infty\) symbol!!

OpenStudy (anonymous):

haa YEAH!

OpenStudy (anonymous):

By the way, \[\sum^\infty_{n=0}{5(-3)^n}=\frac{5}{1-(-3)}=\frac{5}{4}\]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Lol actually I got to go haha so long everybody

OpenStudy (anonymous):

so long...

OpenStudy (tkhunny):

Finding the expression and using the summation notation is your first task. One day, you will be asked to find the sum of expression like these. This is not a really big deal for arithmetic or geometric series. Other series, though, can be a delightful exploration.

OpenStudy (anonymous):

i dont understand what keith did

OpenStudy (tkhunny):

Quite all right. You can add up series later. That is not your assignment for today.

OpenStudy (anonymous):

ok. thank you